# ZoFia: Zero-Shot Fake News Detection with Entity-Guided Retrieval and Multi-LLM Interaction

### v1

@@ 4.1.2 Domain-Specific Expert
It operates in a dynamic and adaptive manner. The system first identifies the most relevant domain from the news text and assigns the agent a precise expert role, such as “economist” or “journalist”.
The expert with this role then analyses along the following 2 dimensions:
@@
This agent uses 3 quadrants of Multi-source Information Matrix: in-news information, LLM internal knowledge, and LLM external knowledge from Wikipedia.
@@ 4.1.3 Claim Verification
Existing research Niu et al. (2024) demonstrates that claim-based fact-checking effectively serves as a reference for LLM-based detection. In our system, a serial pipeline composed of Claim Extractor and Claim Verifier deconstructs and verifies factual claims in the news text. These 2 agents use 3 quadrants of Multi-source Information Matrix: LLM internal knowledge, in-news information, and out-of-news information from Open Web.
Claim Extractor. $\mathcal{M}_{\text{extractor}}$ agent converts unstructured news text $T$ into a set of verifiable structured claims. Its function is formalized as follows:
@@
Claim Verifier. This agent verifies each claim $q$ independently. To ensure precision and control cost Purwar et al. (2024), a simple retrieval augmented generation (RAG) Lewis et al. (2020) is implemented to build a highly relevant context $\mathcal{C}_{\text{rel}}(q)$ for each $q$ from the Open Web corpus $\mathcal{I}_{\text{web}}$.
@@
All decisions are strictly based on $\mathcal{C}_{\text{rel}}(q)$. The final output includes a clear label (“Supports”, “Refutes”, or “Not Enough Information”), a brief reasoning, and evidence directly quoted from $\mathcal{C}_{\text{rel}}(q)$ for suppressing hallucinations.
@@ 5.1 Experimental Setting
Herein we discuss main experimental settings.
Datasets. Following previous state-of-the-art work Liu et al. (2024), our experiments are conducted on two widely recognized fake news datasets: GossipCop and PolitiFact Shu et al. (2020). GossipCop focuses on entertainment, mainly Hollywood celebrity news; PolitiFact focuses on politics, drawing on fact-checks of U.S. political figures. Considering that some links in the original datasets have become invalid, we adopt subsets publicly released in Su et al. (2023).
Metrics and Baselines. We use the accuracy (Acc.) and F1-score (F1) as the main metrics. To reflect performance on detecting fake news, we compute the F1-Score for the fake class. ZoFia compares against two groups of baselines. The first group includes 5 advanced Few-Shot methods: MDFEND Nan et al. (2021), ARG Hu et al. (2024b), PSM Ni et al. (2020), DKFND Liu et al. (2024), and KPL Jiang et al. (2022), with metrics retrieved from the existing literature Jin et al. (2024); Hu et al. (2024c); Liu et al. (2024). The second group consists of 3 representative Zero-Shot methods that we implement for comparison: Standard Prompt, Auto-CoT Zhang et al. (2022), and ReAct Yao et al. (2023) equipped with a search API.
Implementation details. We set the dynamic threshold for entity extraction as $\lambda_{\text{init}}=0.8$ and $\Delta\lambda=0.1$ and the decay factor for MMR as $\gamma=0.5$. The maximum entries for Open Web retrieval is 10; for Wikipedia, we retrieve the first 3 sentences for each entry. The minimum similarity threshold for claim extraction is $\theta_{\text{sim}}=0.1$. The NER model is selected as dslim/bert-base-NER Tjong Kim Sang and De Meulder (2003), and the SBERT model is selected as all-MiniLM-L6-v2 Reimers and Gurevych (2019). All zero-shot methods we implement use DeepSeek-V3 DeepSeek‑AI (2024) with temperature set to 0.1 as the base model, and Brave API [^1] https://brave.com/search/api/ as the search API.
@@ 5.2 Main Experiment
The experimental results are presented in Table 1. ZoFia demonstrates exceptional performance on both datasets, consistently outperforming existing zero-shot and few-shot baselines. On the fact-intensive PolitiFact dataset, ZoFia’s advantages are particularly pronounced, with its accuracy (91.9%) and F1-score (89.8%) not only far exceeding other zero-shot methods but also surpassing all few-shot baselines by a substantial performance margin, proving its powerful efficacy in fact-checking tasks.
Method PolitiFact GossipCop Acc. F1 Acc. F1 MDFEND 65.5 62.3 41.2 40.2 ARG 74.0 67.1 61.4 42.3 PSM 70.0 49.1 77.4 41.7 DKFND 87.0 67.1 82.4 42.3 KPL 58.3 60.4 42.7 42.1 Standard Prompt 79.0 40.2 61.0 28.1 Auto-CoT 89.7 73.7 56.7 43.3 ReAct (Search API) 74.7 67.6 74.0 47.3 ZoFia (Ours) 91.9 89.8 78.5 61.2
The detection task on the GossipCop dataset presents greater challenges, as its content often proves difficult to verify due to subjectivity and factual ambiguity. In this scenario, ZoFia achieves the highest F1-score (61.2%), surpassing all baseline models. Although DKFND attains slightly higher accuracy, its F1-score is notably lower. In contrast, ZoFia’s superior F1-score demonstrates more balanced and robust detection performance across both true and fake news categories while maintaining high precision.
The experimental results clearly indicate that as a zero-shot method, ZoFia consistently outperforms the vast majority of few-shot and zero-shot models, showcasing its advancement and practical utility.
@@ 5.3 Ablation Study
To systematically assess the effectiveness of each component in ZoFia, we conduct extensive ablation on the GossipCop dataset.
Figure 3: Performance comparison of ZoFia’s keyword extraction with other extraction methods.
Effectiveness of Entity-Guided Retrieval. As shown in Figure 3, replacing our keyword extraction module with the classic TF-IDF Sparck Jones (1972) algorithm lowers the F1-Score by 9.1%. This finding emphasizes the need for our Hierarchical Salience. We also evaluate Sentence Salience Bullough et al. (2024) strategy, which retrieves a full sentence. The result shows an extreme imbalance: high Accuracy with low F1-Score. The reason is that using sentences as queries tends to retrieve pages that are highly similar to the original news. This source contamination Deng et al. (2023) induces confirmation bias in LLM agents.
Table 2: Results of ablation study on GossipCop
Component Acc. F1 ZoFia 78.5 61.2 w/o Open Web retrieval 71.7 58.1 w/o Wikipedia retrieval 77.7 59.9 w/o Open Web and Wikipedia retrieval 71.3 58.2 w/o linguist analysis 77.8 58.9 w/o expert analysis 76.8 57.3 w/o linguist and expert analyses 73.5 58.9 w/o claim verification 79.2 60.5 w/o debate 75.0 59.2
Component Analysis of Multi-LLM Interaction. As shown in Table 2, removing Open Web retrieval causes a marked decline, which demonstrates the importance of timely external knowledge. Performance decreases further when all external knowledge sources are removed. For LLM collaboration, removing the linguist or expert weakens the framework. The expert contributes more, with a 3.9% drop in F1-Score. Moreover, removing both causes cumulative degradation. Removing the debate reduces the F1-Score by 2.0%, demonstrating its role in suppressing single-view bias.
@@ 5.4 Sensitivity Analysis of SC-MMR
The dynamic weight $\lambda_{k}$ is the most important parameter of SC-MMR introduced in Section 3. We conduct a sensitivity analysis before designing it. First, $\lambda_{k}$ is fixed to 0.5, indicating equal emphasis on relevance and diversity in keyword selection. As shown in Figure 4, we observe that when the number of selected keywords $k\leq 6$, F1-Score remains stable at a relatively high level, but once $k>6$, F1-Score drops sharply. This pattern indicates that too many keywords introduce noise and hurt detection performance.
@@
Based on this finding, we designed $\lambda_{k}$ a decreasing function with an increasing decay rate, as shown in Eq. 5. It remains high for small $k$ to prioritize relevance, reaching approximately 0.45 at $k=6$, but rapidly decaying for $k>6$. This design lowers the MMR score and smoothly applies the relative termination rule. It allows SC-MMR to adapt the size of the keyword set while avoiding performance loss due to redundancy.
@@ 6 Conclusion
We propose ZoFia, a training-free zero-shot framework for fake news detection, to address the fundamental conflict between the static prior knowledge of large language models (LLMs) and the uncertainty of dynamic news streams. ZoFia uses a two-stage pipeline to tackle this challenge. In the first stage, our novel Hierarchical Salience and SC-MMR algorithms extract a set of informative and low-redundancy keywords from the news text, and then build a multi-source information matrix that integrates LLM internal and external knowledge with in-news and out-of-news information. In the second stage, a dual-state multi-LLM interaction processes this matrix. The agents first perform collaborative analysis to reduce inferential variance, and then conduct adversarial debate to suppress systemic bias, finally delivering an interpretable and robust judgment. Comprehensive experiments on two widely used datasets show that ZoFia outperforms existing zero-shot baselines and the majority of few-shot methods. Detailed ablation study further demonstrates the importance and effectiveness of each core component in the framework.
@@
Though ZoFia show strong detection capacities, its application and evaluation face multiple constraints. Due to copyright and privacy concerns, there has been a recent lack of high-quality, and continuously updated public datasets. It prevents us from evaluating ZoFia on the most recent news. Building the next generation of benchmark datasets that meet ethical standards and reflect real-world information dynamics is a critical step in this field.
From the framework’s perspective, ZoFia depends on multiple LLMs, which introduces nontrivial computation cost and inference latency. The efficiency of using external retrieval can also be improved. Since it is not our main focus, we integrate only a lightweight retrieval augmented generation (RAG) module in claim verification. Future work that adopts more advanced RAG architectures, such as re-ranking model and more advanced mechanism, may further strengthen the exploitation of external knowledge. We plan to conduct more comprehensive experimental benchmark evaluation over the detection capabilities of LLMs and Multi-agent system for fake news Guo et al. (2025); Kuntur et al. (2024); Jiang et al. (2024).

### v3

@@ 4.1.2 Domain-Specific Expert
It operates in a dynamic and adaptive manner. The system first identifies the most relevant domain from the news text and assigns the agent a precise expert role, such as "economist" or "journalist".
The expert with this role then analyzes along the following 2 dimensions:
@@
It uses 3 quadrants of the Multi-Source Information Matrix: in-news information, LLM internal knowledge, and LLM external knowledge.
@@ 4.1.3 Claim Verification
Existing research Niu et al. (2024) demonstrates that claim-based fact checking effectively serves as a reference for LLM-based detection. In our system, a serial pipeline composed of Claim Extractor and Claim Verifier deconstructs and verifies factual claims in the news text. These 2 agents use 3 quadrants of the Multi-Source Information Matrix: LLM internal knowledge, in-news information, and out-of-news information from Open Web.
Claim Extractor. The $\mathcal{M}_{\text{extractor}}$ agent converts unstructured news text $T$ into a set of verifiable structured claims. Its function is formalized as follows:
@@
Claim Verifier. This agent verifies each claim $q$ independently. To ensure precision and control costs Purwar and others (2024), a simple retrieval-augmented generation (RAG) Lewis et al. (2020) is implemented to build a highly relevant context $\mathcal{C}_{\text{rel}}(q)$ for each $q$ from the Open Web corpus $\mathcal{I}_{\text{web}}$.
@@
All decisions are strictly based on $\mathcal{C}_{\text{rel}}(q)$. The final output includes a clear label ("Supports", "Refutes", or "Not Enough Information"), a brief reasoning, and evidence directly quoted from $\mathcal{C}_{\text{rel}}(q)$ to suppress hallucinations.
@@ 4.2 LLM Debate and Judgment
Category Method LLM Usage PolitiFact GossipCop Accuracy F1-Score Accuracy F1-Score Few-shot PSM Nan et al. (2021) Non-LLM 70.00 49.15 77.44 41.73 MDFEND Nan et al. (2021) Non-LLM 65.50 62.30 41.27 40.20 KPL Jiang et al. (2022) Non-LLM 58.33 60.40 42.71 42.08 ARG Hu et al. (2024b) LLM-assisted 74.00 67.16 61.41 42.32 DKFND Liu et al. (2024) LLM-assisted 87.00 82.43 82.37 55.22 Zero-shot Auto-CoT Zhang et al. (2022) LLM-based 89.65 73.67 60.01 48.15 ReAct (Search API) Yao et al. (2023) LLM-based 74.73 67.64 74.03 47.30 HiSS Zhang and Gao (2023) LLM-based 64.82 56.80 68.81 40.40 Web Retrieval Agents Tian et al. (2024) LLM-based 77.83 64.88 66.62 46.54 FactAgent Li et al. (2024b) LLM-based 80.59 70.06 73.80 56.22 CAPE-FND Jin et al. (2025) LLM-based 76.81 62.43 63.79 49.17 DeepSeek-v3 DeepSeek‑AI (2024) Only-LLM 78.99 40.24 61.03 28.17 GPT-4o-mini OpenAI (2024) Only-LLM 73.58 41.66 66.73 33.18 Qwen3-32B Team (2025) Only-LLM 76.57 43.35 63.82 32.45 ZoFia (DeepSeek-v3) LLM-based 91.52 87.88 79.04 61.22 ZoFia (GPT-4o-mini) LLM-based 75.28 75.19 68.90 56.20 ZoFia (Qwen3-32B) LLM-based 84.06 81.76 69.71 59.07
@@ 5.1 Experimental Setting
Datasets. Following previous state-of-the-art work Liu et al. (2024), our experiments are conducted on two widely recognized fake news datasets: GossipCop and PolitiFact Shu et al. (2020). GossipCop focuses on entertainment, mainly Hollywood celebrity news; PolitiFact focuses on politics, drawing on fact-checks of U.S. political figures. Considering that some links in the initial dataset have become invalid, we adopt the available version publicly re-released in Su et al. (2023).
Metrics. We use accuracy and macro F1-Score as evaluation metrics. F1-Score is less affected by data imbalance, so it serves as the primary metric for assessment.
Baselines. We incorporate two groups of baselines. The first group includes advanced few-shot methods: PSM Ni et al. (2020), MDFEND Nan et al. (2021), ARG Hu et al. (2024b), DKFND Liu et al. (2024), and KPL Jiang et al. (2022), with reliable metrics from the existing works Jin et al. (2024); Hu et al. (2024c); Liu et al. (2024). The second group consists of representative zero-shot methods for comparison: Auto-CoT Zhang et al. (2022), ReAct Yao et al. (2023) equipped with a search API, HiSS Zhang and Gao (2023), Web Retrieval Agents Tian et al. (2024), FactAgent Li et al. (2024b), and CAPE-FND Jin et al. (2025).
We implement ZoFia based on DeepSeek-v3 DeepSeek‑AI (2024), GPT-4o-mini OpenAI (2024) and Qwen3-32B Team (2025), and compare it with only-LLM inference. All other LLM-based zero-shot methods are based on DeepSeek-v3.
Implementation details. We set the dynamic threshold of entity extraction as $\lambda_{\text{init}}=0.8$ and $\Delta\lambda=0.1$ and the decay factor of MMR as $\gamma=0.5$. The maximum number of entries for Open Web retrieval is 10; for Wikipedia, we retrieve the first 3 sentences for each entry. The minimum similarity threshold for claim extraction is $\theta_{\text{sim}}=0.1$. The NER model is selected as dslim/bert-base-NER Tjong Kim Sang and De Meulder (2003), and the SBERT model is selected as all-MiniLM-L6-v2 Reimers and Gurevych (2019). For all base LLMs, we set the temperature to 0.
Brave API [^1] https://brave.com/search/api/ is selected as Open Web retrieval API. To strictly prevent label information leakage, we set the search cutoff date to the day before the publication date of each sample URL and exclude dataset-associated fact-checking domains from open-web retrieval (gossipcop.com, politifact.com, and snopes.com).
@@ 5.2 Main Experiment
ZoFia demonstrates exceptional performance on both datasets as shown in Table 1, consistently outperforming existing zero-shot and few-shot baselines. On the fact-intensive PolitiFact dataset, ZoFia’s advantages are particularly pronounced, with its accuracy (91.52%) and F1-Score (87.88%) not only far exceeding other zero-shot methods but also surpassing all few-shot baselines by a substantial performance margin.
The detection task on the GossipCop dataset presents greater challenges, as its content often proves difficult to verify due to subjectivity and factual ambiguity. In this scenario, ZoFia achieves the highest F1-Score (61.22%), surpassing all baseline models. Although DKFND attains slightly higher accuracy, its F1-Score is notably lower. In contrast, ZoFia’s superior F1-Score demonstrates more balanced and robust detection performance across both true and fake news categories while maintaining high precision.
The results clearly show that as a zero-shot method, ZoFia consistently outperforms all zero-shot baselines and even most few-shot methods. Compared to only-LLM inference, ZoFia provides stable and substantial performance gains, demonstrating the superiority of its architecture.
@@ 5.3 Ablation Study
To assess the contribution of each component in ZoFia, we conduct ablation studies on GossipCop with DeepSeek-v3 and Qwen3-32B. As shown in Table 2, removing Open Web retrieval causes clear degradation in both settings, which shows that timely external evidence is essential for both base models. Removing all retrieval sources further lowers F1-Score to 58.26% on DeepSeek-v3 and 50.85% on Qwen3-32B. In LLM collaboration, removing either the linguist or the expert weakens the framework, and the expert contributes more; removing both leads to a larger cumulative drop. Removing debate also consistently reduces F1-Score, which shows that it helps mitigate single-perspective bias. The F1-Score drops by 2.06% on DeepSeek-v3 and by 5.00% on Qwen3-32B, and this larger drop suggests that the smaller model is more sensitive to the loss of this mechanism.
Table 2: Ablation study results on GossipCop. (ACC.: Accuracy, F1: F1-Score)
Component DeepSeek-v3 Qwen3-32B Acc. F1 Acc. F1 ZoFia 79.04 61.22 69.71 59.07 w/o Wikipedia retrieval 77.67 59.90 68.09 56.58 w/o Open Web retrieval 71.71 58.14 64.18 52.47 w/o All retrievals 71.37 58.26 62.65 50.85 w/o Linguist analysis 77.86 58.87 67.55 54.62 w/o Expert analysis 76.77 57.32 66.87 54.39 w/o All analyses 73.51 57.90 63.82 52.71 w/o Claim verification 79.23 60.50 67.48 55.17 w/o Debate 75.01 59.16 67.13 54.07
5.4 Effectiveness of Entity-Guided Retrieval
To verify the effectiveness of entity-guided retrieval in ZoFia, we fix the second stage and vary only how the query is constructed on GossipCop with Qwen3-32B. TF-IDF Sparck Jones (1972), PromptNER Ashok and Lipton (2023), and Only-LLM NER still build queries from keywords, but replace the entity acquisition mechanism, and their F1-Scores drop by 5.69%, 3.94%, and 6.33% relative to ZoFia. Sentence Salience Bullough et al. (2024), Core Claim, and OR Composition instead use sentence-level queries or relax conjunctive matching from AND ($\wedge$) to OR ($\vee$), and their F1-Scores fall by 12.92%, 10.26%, and 6.81%. These results show that the gain does not come from arbitrary keyword lists, and it depends on accurate entity extraction and strict entity-level constraints, which make the original conjunctive entity-guided query the most effective design.
Table 3: Effectiveness of query construction variants in ZoFia on GossipCop.
Method Accuracy F1-Score
ZoFia 69.71 59.07
ZoFia (w/o All retrievals) 62.65 50.85
ZoFia (TF-IDF) 66.97 53.38
ZoFia (PromptNER) 63.08 55.13
ZoFia (Only-LLM NER) 61.69 52.74
ZoFia (Sentence Salience) 68.64 46.15
ZoFia (Core Claim) 66.65 48.81
ZoFia (OR Composition) 64.89 52.26
@@ 5.5 Sensitivity Analysis of SC-MMR
To motivate the design of the dynamic weight $\lambda_{k}$, we first conduct an experiment with the weight fixed at $\lambda=0.5$ to observe the direct impact of the keyword count $k$ on performance. As shown in Figure 3, the F1-Score remains stable when $k\leq 6$ but drops sharply beyond this point.
@@
This phenomenon reveals an inflection point: when the number of keywords exceeds $6$, the risk of introducing noise and redundancy begins to outweigh the benefits of new information. It suggests that the strategic focus must shift from relevance to diversity near the critical inflection point of $k\approx 6$. Consequently, we adopt the annealing schedule form defined in Equation 5 and determine its parameters as $\alpha=0.3$ and $\beta=2.5$ to fit this downward trend.
5.6 Efficiency of Token Utilization
To investigate ZoFia’s reasoning efficiency, a controlled comparison experiment is conducted. We provide CoT Wei et al. (2022), FactAgent Li et al. (2024b), and ZoFia with the same materials and impose a unified limit on output tokens to evaluate their performance. Since ZoFia cannot complete full reasoning below 400 tokens, we set its evaluation range to 400 tokens and above.
Figure 4: Performance (F1-Score) comparison of methods under maximum output token limits.
The results are shown in Figure 4. At all budget points above 400 tokens, ZoFia achieves a clearly higher F1-Score than other baseline methods, which demonstrates high reasoning efficiency. In contrast to CoT and FactAgent, whose performance saturates after around 600 tokens, ZoFia shows a stable upward trend even at 1200 tokens. This indicates that ZoFia can consistently convert tokens into performance gains within an acceptable budget.
Note that this experiment only constrains output tokens, because prompt caching Gim et al. (2024) can extensively amortize input overhead, making output cost more critical for efficiency.
@@ 6 Conclusion
In this paper, we propose ZoFia, a retrieval-augmented multi-agent zero-shot framework for fake news detection, to address the cognitive conflict of a single LLM between content reasoning and fact checking. It first utilizes our novel Hierarchical Salience and Salience-Calibrated Minimum Marginal Relevance (SC-MMR) algorithm to accurately extract core entities from news text, which then guide dual-source retrieval from Open Web and Wikipedia. Next, the multi-agent system conducts analysis and verification in parallel and makes a final verdict via adversarial debate. This process effectively reduces confirmation bias from a single reasoning perspective and ensures robust and explainable results. Comprehensive experiments on two public datasets show that ZoFia outperforms existing zero-shot baselines in both performance and efficiency.
@@
Though ZoFia shows strong detection capabilities, its application and evaluation face multiple constraints. Due to copyright and privacy concerns, there has been a recent lack of high-quality, continuously updated public datasets. It prevents us from evaluating ZoFia on the most recent news. Building the next generation of benchmark datasets that meet ethical standards and reflect real-world information dynamics is a critical step in this field.
The efficiency of using external retrieval can also be improved. Since it is not our main focus, we integrate only a lightweight retrieval-augmented generation (RAG) module for claim verification. Future work that adopts more advanced RAG architectures, such as a re-ranking model and more advanced mechanisms, may further strengthen the exploitation of external knowledge. We plan to conduct more comprehensive benchmark evaluations of the detection capabilities of LLMs and multi-agent systems for fake news Guo et al. (2025); Kuntur et al. (2024); Jiang et al. (2024).